# Computable Scott sentences and the weak Whitehead problem for finitely presented groups

### 29 Dec 2023 v1

@@ 1. Introduction
In [21] Scott proved that every countable structure is characterized up to isomorphism by a sentence of $\mathfrak{L}_{\omega_{1},\omega}$. Recently, there has been considerable interest [2, 6, 8, 9, 10, 11, 15] in understanding how simple such a sentence can been chosen, in the sense of following stratification of $\mathfrak{L}_{\omega_{1},\omega}$ from computable model theory:
@@
For a general overview of the motivation, history and literature surrounding the topic of optimal Scott sentences we refer the reader to the excellent survey [8]. In this paper we deal exclusively with optimal (computable) Scott sentences for finitely generated structures, and in particular for finitely presented groups; this topic has been central to the area. We shortly review the main advancements in this direction. In [15] it was proved that every finitely generated structure has a $\Sigma_{3}$ Scott sentence. In the same paper it was proved that several familiar finitely generated structures (e.g. free groups of finite rank, free abelian groups of finite rank, the infinite dihedral group) actually have a $d$-$\Sigma_{2}$ Scott sentence, and in fact a computable one. This led the authors to wonder if indeed every finitely generated group has a $d$-$\Sigma_{2}$ Scott sentence. This was answered negatively in [10] using methods from combinatorial group theory. In [20] we isolated some general sufficient conditions for the existence of a computable $d$-$\Sigma_{2}$ Scott sentence, thus deducing e.g. that right-angled Coxeter groups of finite rank have a computable $d$-$\Sigma_{2}$ Scott sentence. Many questions on optimal Scott sentences for finitely generated structures are still open (see in particular [8, 9] for an overview). The following two questions appear to be the most important in the area (cf. [9, Questions 1 and 3] and references therein):
@@
For the rest of this introduction, we will shorten finitely presented as f.p. Concerning the first question, in [9] Harrison-Trainor gave a purely group theoretic characterization of the f.p. groups without a $d$-$\Sigma_{2}$ Scott sentence; despite this the question remains wide open at the moment. Our paper deals instead with Question 1.2. Concerning the existence of computable Scott sentences for f.p. (or more generally finitely generated) groups little appears to be known (cf. [8, pg. 12]). An important conceptual result is in [9], in this paper Harrison-Trainor constructs a finitely generated module with a $d$-$\Sigma_{2}$ Scott sentence but no computable $d$-$\Sigma_{2}$ Scott sentence. From this, using the universality among finitely generated structures of finitely generated groups [10], Harrison-Trainor infers the abstract existence of a finitely generated group with a $d$-$\Sigma_{2}$ Scott sentence but no computable $d$-$\Sigma_{2}$ Scott sentence, but no explicit example of such a group can be exhibited using his argument. Furthermore, the argument of Harrison-Trainor from [9] says nothing on finite presentability. The only known general results that we are aware of come from our paper [20], already mentioned at the introduction, but the condition isolated there, although useful for some concrete applications, appeared to be of a rather technical nature. In this paper we show that, under the assumption of Hopfianity, the existence of a computable $d$-$\Sigma_{2}$ Scott sentence is equivalent to a weakening of a well-known algorithmic problem from combinatorial group theory: the Whitehead problem. This problem asks about the decidability of determining whether two tuples in a f.p. group are automorphic. The problem was first solved for free groups by Whitehead in [26] (hence the name). There are now several important results of this kind, see in particular references [4, 13]. We define the weak Whithead problem.
@@
Let $A$ be a computable Hopfian f.p. structure, then $A$ has a computable $d$-$\Sigma_{2}$ Scott sentence iff the weak Whitehead problem for $A$ is decidable.
@@
There exists an Hopfian finitely presented group with decidable word problem (so a computable such group) but undecidable weak Whithead problem.
@@ 2. The proof of the characterization
Given a finitely generated structure $A=\langle\bar{a}\rangle$ we say that $(A,\bar{a})$ has decidable weak Whitehead problem if there exists an algorithm which decides whether a tuple $\bar{b}\in G^{n}$ is automorphic to $\bar{a}$ (i.e., if there is $\alpha\in\mathrm{Aut}(A)$ such that $\alpha(\bar{a})=\bar{b}$).
@@
The fact that (2) implies (1) is by Lemma 2.9. Concerning the other direction, let $\Delta(\bar{x})$ be a computable $\Pi_{1}$-formula definining the $\mathrm{Aut}(A)$-orbit of $\bar{a}$. Evidently, deciding the weak Whitehead problem for $A=\langle\bar{a}\rangle$ means deciding (algorithmically) whether a tuple $\bar{b}\in A^{n}$ is or not in the following set:
@@ 1. $(\star_{1})$
Why $(\star_{1})$? Recall that $\Delta(\bar{x})$ is a computable $\Pi_{1}$-formula definining the $\mathrm{Aut}(A)$-orbit of $\bar{a}$ and so $\bar{b}\in X_{*}$ if and only if $A\models\Delta(\bar{b})$. It follows that $\bar{b}\in\widehat{X}_{*}$ iff $A\models\neg\Delta(\bar{b})$, but the negation of a $\Pi_{1}$-formula is a $\Sigma_{1}$-formula, and by [3, Theorem 7.5], in a computable structure a $\Sigma_{1}$-formula always defines a computably enumerable set.
@@ 1. $(\star_{2})$
We recall the definition of $A$ having decidable weak Whitehead problem.
@@ 3. Groups defined by their types
Let $\mathcal{F}$ be a set of first-order formulas in the variables $\{x_{i}:i<\omega\}$ such that $\mathcal{F}$ contains all the positive quantifier-free formulas in the group theory language, $\mathcal{F}$ is closed under $\wedge$ and $\vee$, and $\mathcal{F}$ is closed under substitution of variables in words in $\{x_{i}:i<\omega\}$. We say that $f:A\rightarrow B$ is an $\mathcal{F}$-embedding if for every $\varphi(\bar{a})\in\mathcal{F}$ and $\bar{a}\in A^{\mathrm{lg}(\bar{x})}$ we have that $A\models\varphi(\bar{x})$ if and only if $B\models\varphi(f(\bar{a}))$. We say that a finitely generated group $A$ (or, more generally, structure) is strongly defined by its $\mathcal{F}$-types if and only if any $\mathcal{F}$-embedding $f:A\rightarrow A$ is an automorphism of $A$.
@@
Is suffices to show that hyperbolic groups are finitely presented and Hopfian. The fact that they are finitely presented is well-known. Concerning Hopfianity, see e.g. [24]. ∎

### 04 Mar 2024 v2

@@ 1. Introduction
In [21] Scott proved that every countable structure is characterized up to isomorphism by a sentence in $\mathfrak{L}_{\omega_{1},\omega}$. Recently, there has been considerable interest [2, 6, 8, 9, 10, 11, 15] in understanding how simple such a sentence can been chosen, in the sense of following stratification of $\mathfrak{L}_{\omega_{1},\omega}$ from computable model theory:
@@
For a general overview of the motivation, history and literature surrounding the topic of optimal Scott sentences we refer the reader to the excellent survey [8]. In this paper we deal exclusively with optimal (computable) Scott sentences for finitely generated structures, and in particular for finitely presented groups; this topic has been central to the area. We shortly review the main progress in this direction. In [15] it was proved that every finitely generated structure has a $\Sigma_{3}$ Scott sentence. In the same paper it was proved that several familiar finitely generated structures (e.g. free groups of finite rank, free abelian groups of finite rank, the infinite dihedral group) actually have a $d$-$\Sigma_{2}$ Scott sentence, and in fact a computable one. This led the authors to wonder if indeed every finitely generated group has a $d$-$\Sigma_{2}$ Scott sentence. This was answered negatively in [10] using methods from combinatorial group theory. In [20] we isolated some general sufficient conditions for the existence of a computable $d$-$\Sigma_{2}$ Scott sentence, thus deducing e.g. that right-angled Coxeter groups of finite rank have a computable $d$-$\Sigma_{2}$ Scott sentence. Many questions on optimal Scott sentences for finitely generated structures are still open (see in particular [8, 9] for an overview). The following two questions appear to be the most important in the area (cf. [9, Questions 1 and 3] and references therein):
@@
For the rest of this introduction, we will shorten finitely presented as f.p. Concerning the first question, in [9] Harrison-Trainor gave a purely group theoretic characterization of the f.p. groups without a $d$-$\Sigma_{2}$ Scott sentence; despite this the question remains wide open at the moment. Our paper deals instead with Question 1.2. Concerning the existence of computable Scott sentences for f.p. (or more generally finitely generated) groups little appears to be known (cf. [8, pg. 12]). An important conceptual result is in [9], in this paper Harrison-Trainor constructs a finitely generated module with a $d$-$\Sigma_{2}$ Scott sentence but no computable $d$-$\Sigma_{2}$ Scott sentence. From this, using the universality among finitely generated structures of finitely generated groups [10], Harrison-Trainor infers the abstract existence of a finitely generated group with a $d$-$\Sigma_{2}$ Scott sentence but no computable $d$-$\Sigma_{2}$ Scott sentence, but no explicit example of such a group can be exhibited using his argument. Furthermore, the argument of Harrison-Trainor from [9] says nothing on finite presentability. The only known general results that we are aware of come from our paper [20], already mentioned at the introduction, but the condition isolated there, although useful for some concrete applications, appeared to be of a rather technical nature. In this paper we show that, under the assumption of Hopfianity, the existence of a computable $d$-$\Sigma_{2}$ Scott sentence is equivalent to a weakening of a well-known algorithmic problem from combinatorial group theory: the Whitehead problem. This problem asks about the decidability of determining whether two tuples in a f.p. group are automorphic. The problem was first solved for free groups by Whitehead in [26] (hence the name). There are now several important results of this kind, see in particular references [4, 13]. We define the weak Whitehead problem.
@@
Let $A$ be a computable Hopfian f.p. structure. Then $A$ has a computable $d$-$\Sigma_{2}$ Scott sentence iff the weak Whitehead problem for $A$ is decidable.
@@
There exists a Hopfian finitely presented group with decidable word problem (so a computable such group) but undecidable weak Whitehead problem.
@@ 2. The proof of the characterization
Given a finitely generated structure $A=\langle\bar{a}\rangle$ we say that $(A,\bar{a})$ has decidable weak Whitehead problem if there exists an algorithm which decides whether a tuple $\bar{b}\in G^{n}$ is automorphic to $\bar{a}$ (i.e., if $\exists\alpha\in\mathrm{Aut}(A)$ s.t. $\alpha(\bar{a})=\bar{b}$).
@@
The fact that (2) implies (1) is by Lemma 2.9. Concerning the other direction, let $\Delta(\bar{x})$ be a computable $\Pi_{1}$-formula defining the $\mathrm{Aut}(A)$-orbit of $\bar{a}$. Evidently, deciding the weak Whitehead problem for $A=\langle\bar{a}\rangle$ means deciding (algorithmically) whether a tuple $\bar{b}\in A^{n}$ is or not in the following set:
@@ 1. $(\star_{1})$
Why $(\star_{1})$? Recall that $\Delta(\bar{x})$ is a computable $\Pi_{1}$-formula defining the $\mathrm{Aut}(A)$-orbit of $\bar{a}$ and so $\bar{b}\in X_{*}$ if and only if $A\models\Delta(\bar{b})$. It follows that $\bar{b}\in\widehat{X}_{*}$ iff $A\models\neg\Delta(\bar{b})$, but the negation of a $\Pi_{1}$-formula is a $\Sigma_{1}$-formula, and by [3, Theorem 7.5], in a computable structure a $\Sigma_{1}$-formula always defines a computably enumerable set.
@@ 1. $(\star_{2})$
We recall the definition of $A$ having a decidable weak Whitehead problem.
@@ 3. Groups defined by their types
Let $\mathcal{F}$ be a set of first-order formulas in the variables $\{x_{i}:i<\omega\}$ such that $\mathcal{F}$ contains all the positive quantifier-free formulas in the group theory language, $\mathcal{F}$ is closed under $\wedge$ and $\vee$, and $\mathcal{F}$ is closed under substitution of variables in words in $\{x_{i}:i<\omega\}$. We say that $f:A\rightarrow B$ is an $\mathcal{F}$-embedding if, for every $\varphi(\bar{a})\in\mathcal{F}$ and $\bar{a}\in A^{\mathrm{lg}(\bar{x})}$, we have that $A\models\varphi(\bar{x})$ if and only if $B\models\varphi(f(\bar{a}))$. We say that a finitely generated group $A$ (or, more generally, structure) is strongly defined by its $\mathcal{F}$-types iff any $\mathcal{F}$-embedding $f:A\rightarrow A$ is an automorphism of $A$.
@@
It suffices to show that hyperbolic groups are finitely presented and Hopfian. The fact that they are finitely presented is well-known. Concerning Hopfianity, see e.g. [24]. ∎